# SPAN: Spatial-Projection Alignment for Monocular 3D Object Detection

### v1

@@ Abstract
Existing monocular 3D detectors typically tame the pronounced nonlinear regression of 3D bounding box through decoupled prediction paradigm, which employs multiple branches to estimate geometric center, depth, dimensions, and rotation angle separately. Although this decoupling strategy simplifies the learning process, it inherently ignores the geometric collaborative constraints between different attributes, resulting in the lack of geometric consistency prior, thereby leading to suboptimal performance. To address this issue, we propose novel Spatial-Projection Alignment (SPAN) with two pivotal components: (i). Spatial Point Alignment enforces an explicit global spatial constraint between the predicted and ground‑truth 3D bounding boxes, thereby rectifying spatial drift caused by decoupled attribute regression. (ii). 3D-2D Projection Alignment ensures that the projected 3D box is aligned tightly within its corresponding 2D detection bounding box on the image plane, mitigating projection misalignment overlooked in previous works. To ensure training stability, we further introduce a Hierarchical Task Learning strategy that progressively incorporates spatial-projection alignment as 3D attribute predictions refine, preventing early stage error propagation across attributes. Extensive experiments demonstrate that the proposed method can be easily integrated into any established monocular 3D detector and delivers significant performance improvements.
@@ 1 Introduction
Our work addresses this gap by explicitly integrating Spatial-Projection Alignment into an end-to-end framework. To be concrete as shown in Fig. 1, our central idea is to impose geometrically collaborative constraints on every set of 3D bounding box attributes and to optimize them jointly to enhance both spatial and projection coherence. Specifically, we first propose Spatial Point Alignment. In contrast to previous works ROI-10D [23] and MonoDIS [33] that simply regress corners as an auxiliary task, our loss is applied to the eight corner coordinates derived from the primary 7-DoF attributes. This directly regularizes the main box parameters by constraining them to align with the ground-truth corners, thereby enforcing geometric integrity. Second, we propose 3D-2D Projection Alignment. Unlike analytical solvers like Deep3DBox [24] that are sensitive to 2D detection noise, our differentiable approach provides a robust, gradient-based learning signal. It projects the corners of each predicted 3D box onto the image plane and constrains the minimal enclosing rectangle of this projection to align with its corresponding 2D detection box.
During early training, the predicted 3D bounding box parameters are subject to substantial noise, rendering our geometric collaborative constraints unreliable. Consequently, after introducing the Spatial-Projection Alignment, we observed marked instability at the outset of model training. To mitigate this, we adapted and refined GUPNet’s [20] Hierarchical Task Learning strategy. Briefly, this scheme assigns greater weight to the 2D and 3D attribute regression objectives during the initial training epochs, and then gradually increases the emphasis on the more challenging geometric collaborative constraints as learning progresses, preserving training stability in the early phase.
@@
  • We identify that the prevailing decoupled regression paradigm in monocular 3D detection ignores the inherent spatial and projection relationships among bounding-box attributes. This oversight leads to inconsistent predictions that violate physical projection constraints and ultimately caps localization accuracy.
@@
  • This Spatial-Projection Alignment can be employed as a plug-and-play module within mainstream monocular 3D detectors. Experimental results reveal that it elevates the accuracy of state-of-the-art monocular 3D detectors without introducing additional modules.
@@ 2 Related Work
Monocular 3D Object Detection. Monocular 3D object detection seeks to recover 3D bounding boxes from a single image. However, the absence of direct depth cues renders this an ill-posed problem. Early works such as Deep3DBox [24] and Shift R-CNN [25] leverage 2D to 3D perspective geometry constraints to locate the 3D box center. More recent approaches [20, 18, 14, 44, 8, 43, 41, 29] combine data-driven frameworks with learnable geometric priors. For example, [20, 9, 21] introduce uncertainty into geometric projection-based depth estimation to improve detection accuracy. MonoDETR [44] incorporates an object-level foreground depth map and fuses regressed depth, geometric depth, and foreground depth to obtain a final depth estimate. FD3D [41] further supervises intermediate features using the foreground depth map to provide clearer guidance for foreground objects. MonoCD [43] systematically analyzes the complementarity of diverse depth cues in geometric space and fuses them to mitigate positive and negative errors. MonoDGP [29] studies geometric error in projected depth and proposes a perspective invariant correction formula. Although these methods have achieved notable progress, the produced 3D boxes still exhibit geometric misalignment because each parameter is optimized independently, lacking spatial coherence. Our geometric collaborative loss alleviates this issue by imposing unified geometric constraints during end-to-end training.
@@ 3.2 Spatial Point Alignment
$\begin{gathered}\mathrm{proj}_{\mathbf{P},\mathbf{a}_{k}}=\{\mathbf{P}_{i}\cdot\mathbf{a}_{k}|i=1,\cdots,8\},\\ \mathrm{~proj}_{\mathbf{G},\mathbf{a}_{k}}=\{\mathbf{G}_{i}\cdot\mathbf{a}_{k}|i=1,\cdots,8\}.\end{gathered}$ (3)
@@
$\begin{gathered}\left|{{\mathbf{P}_{{{\mathbf{a}}_{k}}}}}\right|={\max}_{{\mathbf{P}},{{\mathbf{a}}_{k}}}-{{\min}_{{\mathbf{P}},{{\mathbf{a}}_{k}}}},\left|{{\mathbf{G}_{{{\mathbf{a}}_{k}}}}}\right|={{{\max}_{{\mathbf{G}},{{\mathbf{a}}_{k}}}}}-{{\min}_{{\mathbf{G}},{{\mathbf{a}}_{k}}}},\\ \left|\mathbf{C}\right|=\max\left({{{\max}_{{\mathbf{P}},{{\mathbf{a}}_{k}}}},{{\max}_{{\mathbf{G}},{{\mathbf{a}}_{k}}}}}\right)-\min\left({{{\min}_{{\mathbf{P}},{{\mathbf{a}}_{k}}}},{{\min}_{{\mathbf{G}},{{\mathbf{a}}_{k}}}}}\right),\\ \mathrm{GIoU}_{k}^{1D}=\frac{\left|{{\mathbf{P}_{{{\mathbf{a}}_{k}}}}}\cap{{\mathbf{G}_{{{\mathbf{a}}_{k}}}}}\right|}{\left|{{\mathbf{P}_{{{\mathbf{a}}_{k}}}}}\cup{{\mathbf{G}_{{{\mathbf{a}}_{k}}}}}\right|}-\frac{\left|\mathbf{C}\right|-\left|{{\mathbf{P}_{{{\mathbf{a}}_{k}}}}}\cup{{\mathbf{G}_{{{\mathbf{a}}_{k}}}}}\right|}{\left|\mathbf{C}\right|},\\ \mathrm{MGIoU}^{3D}=\frac{1}{3}\sum_{k=1}^{3}\mathrm{GIoU}_{k}^{1D}.\end{gathered}$ (5)
@@ 3.4 Total Loss and Training Strategy
Although SPAN promotes joint optimization of the 3D attributes and markedly improves both spatial and projection consistency, it can destabilize early training. Specifically, noisy initial predictions of the 3D box parameters can invalidate the assumptions underlying these high‑order constraints. To mitigate this, we adopt and adapt the Hierarchical Task Learning (HTL) strategy [20]. HTL applies a time‑varying weight ${\omega_{i}}\left(t\right)$ to each loss term, dynamically adjusting these weights at each epoch based on the learning progress of their prerequisite tasks. The implementation of HTL is provided in the appendix Sec. C.
@@
As illustrated in Fig. 4, the overall learning procedure is organized into four sequential stages. Stage 1 focuses on 2D detection, covering object classification, 2D box localization, and projected center regression. Stage 2 addresses 3D dimension and rotation angle regression. Following MonoDGP [29], whose decoder is decoupled and initializes 3D object queries from 2D cues, we treat the tasks of Stage 1 as pre-tasks for Stage 2. Stage 3 performs depth estimation, since depth relies on the geometric relationship between 2D box height and 3D object height, its prerequisites encompass all tasks from both Stage 1 and Stage 2. Finally, Stage 4 introduces the Spatial-Projection Alignment, which depends on all preceding 3D attribute regression tasks. This staged design ensures that each task commences only after its prerequisite tasks have stabilized, thereby preserving training stability throughout the optimization process.
@@ A Detailed Discussion on Projection Constraint
In summary, the extrema of $f\left({{x_{n}},{z_{n}}}\right)$ over the point set $S$ are necessarily attained in at least one of the four line segments $(x_{\min},y_{n},z_{\min})$, $(x_{\min},y_{n},z_{\max})$, $(x_{\max},y_{n},z_{\min})$, and $(x_{\max}y_{n},z_{\max})$. Since the vertices of the 3D bounding box are precisely the endpoints of these segments, the maximum and minimum values of $u_{n}$ over $S$ are realized at these vertices, the identical argument holds for $v_{n}$.
@@
${u_{\min}}\leqslant{u_{n}}\leqslant{u_{\max}},{v_{\min}}\leqslant{v_{n}}\leqslant{v_{\max}},$ (16)
@@ B Detailed Loss Function
$\mathcal{L}_{region}^{i}=1-\frac{2\cdot\sum_{j=1}^{N_{i}}p_{j}\cdot g_{j}}{\sum_{j=1}^{N_{i}}p_{j}+\sum_{j=1}^{N_{i}}g_{j}},$ (21)
@@ D Experiments on Waymo Open Dataset
E More Qualitative Results

### v2

@@ Abstract
Existing monocular 3D detectors typically tame the pronounced nonlinear regression of 3D bounding box through decoupled prediction paradigm, which employs multiple branches to estimate geometric center, depth, dimensions, and rotation angle separately. Although this decoupling strategy simplifies the learning process, it inherently ignores the geometric collaborative constraints between different attributes, resulting in the lack of geometric consistency prior, thereby leading to suboptimal performance. To address this issue, we propose novel Spatial-Projection Alignment (SPAN) with two pivotal components: (i). Spatial Point Alignment enforces an explicit global spatial constraint between the predicted and ground‑truth 3D bounding boxes, thereby rectifying spatial drift caused by decoupled attribute regression. (ii). 3D-2D Projection Alignment ensures that the projected 3D box is aligned tightly within its corresponding 2D detection bounding box on the image plane, mitigating projection misalignment overlooked in previous works. To ensure training stability, we further introduce a Hierarchical Task Learning strategy that progressively incorporates spatial-projection alignment as 3D attribute predictions refine, preventing early stage error propagation across attributes. Extensive experiments demonstrate that the proposed method can be easily integrated into any established monocular 3D detector and delivers significant performance improvements. Project Page: https://wyfdut.github.io/SPAN/
@@ 1 Introduction
Our work addresses this gap by explicitly integrating Spatial-Projection Alignment into an end-to-end framework. To be concrete as shown in Fig. 1, our central idea is to impose geometrically collaborative constraints on every set of 3D bounding box attributes and to optimize them jointly to enhance both spatial and projection coherence. Specifically, we first propose Spatial Point Alignment. In contrast to ROI-10D [24] and MonoDIS [34] that simply regress corners as an auxiliary task, our loss is applied to the eight corner coordinates derived from the primary 7-DoF attributes. This directly regularizes the main box parameters by constraining them to align with the ground-truth corners, thereby enforcing geometric integrity. Second, we propose 3D-2D Projection Alignment. Unlike analytical solvers like Deep3DBox [25] that are sensitive to 2D detection noise, our differentiable approach provides a robust, gradient-based learning signal. It projects the corners of each predicted 3D box onto the image plane and constrains the minimal enclosing rectangle of this projection to align with its corresponding 2D detection box.
During early training, the predicted 3D bounding box parameters are subject to substantial noise, rendering our geometric collaborative constraints unreliable. Consequently, after introducing the Spatial-Projection Alignment, we observed marked instability at the outset of model training. To mitigate this, we adapt and refine GUPNet’s [22] Hierarchical Task Learning strategy. Briefly, this scheme assigns greater weight to the 2D and 3D attribute regression objectives during the initial training epochs, and gradually increases the emphasis on the more challenging geometric collaborative constraints as learning progresses, preserving training stability in the early phase.
@@
  • We observe that the prevailing decoupled regression paradigm in monocular 3D detection often overlooks the inherent spatial and projection relationships among the bounding-box attributes. This gap can result in predictions that violate physical projection constraints, thereby limiting localization accuracy.
@@
  • The proposed Spatial-Projection Alignment can be employed as a plug-and-play module within mainstream monocular 3D detectors. Experimental results reveal that it elevates the accuracy of state-of-the-art monocular 3D detectors without introducing additional modules.
@@ 2 Related Work
Monocular 3D Object Detection. Monocular 3D object detection seeks to recover 3D bounding boxes from a single image. However, the absence of direct depth cues renders this an ill-posed problem. Early works such as Deep3DBox [25] and Shift R-CNN [26] leverage 2D to 3D perspective geometry constraints to locate the 3D box center. More recent approaches [22, 19, 14, 46, 8, 45, 43, 30] combine data-driven frameworks with learnable geometric priors. For example, [22, 9, 21] introduce uncertainty into geometric projection-based depth estimation to improve detection accuracy. [38] employs a hybrid feature aggregation technique [47, 16] to enhance the receptive field.  MonoDETR [46] incorporates an object-level foreground depth map and fuses regressed depth, geometric depth, and foreground depth to obtain a final depth estimate. FD3D [43] further supervises intermediate features using the foreground depth map to provide clearer guidance for foreground objects. MonoCD [45] systematically analyzes the complementarity of diverse depth cues in geometric space and fuses them to mitigate positive and negative errors. MonoDGP [30] studies geometric error in projected depth and proposes a perspective invariant correction formula. Although these methods have achieved notable progress, the produced 3D boxes still exhibit geometric misalignment because each parameter is optimized independently, lacking spatial coherence. Our geometric collaborative loss alleviates this issue by imposing unified geometric constraints during end-to-end training.
@@ 3.2 Spatial Point Alignment
$\begin{gathered}\mathrm{proj}^{\left(i\right)}_{\mathbf{P},\mathbf{a}_{k}}=\mathbf{P}_{i}\cdot\mathbf{a}_{k},\\ \mathrm{~proj}^{\left(i\right)}_{\mathbf{G},\mathbf{a}_{k}}=\mathbf{G}_{i}\cdot\mathbf{a}_{k}.\end{gathered}$ (3)
@@
$\begin{gathered}{{\mathbf{P}_{{{\mathbf{a}}_{k}}}}}=[{{\min}_{{\mathbf{P}},{{\mathbf{a}}_{k}}}},{\max}_{{\mathbf{P}},{{\mathbf{a}}_{k}}}],{{\mathbf{G}_{{{\mathbf{a}}_{k}}}}}=[{{\min}_{{\mathbf{G}},{{\mathbf{a}}_{k}}}},{{{\max}_{{\mathbf{G}},{{\mathbf{a}}_{k}}}}}],\\ \left|\mathbf{C}\right|=\max\left({{{\max}_{{\mathbf{P}},{{\mathbf{a}}_{k}}}},{{\max}_{{\mathbf{G}},{{\mathbf{a}}_{k}}}}}\right)-\min\left({{{\min}_{{\mathbf{P}},{{\mathbf{a}}_{k}}}},{{\min}_{{\mathbf{G}},{{\mathbf{a}}_{k}}}}}\right),\\ \mathrm{GIoU}_{k}^{1D}=\frac{\left|{{\mathbf{P}_{{{\mathbf{a}}_{k}}}}}\cap{{\mathbf{G}_{{{\mathbf{a}}_{k}}}}}\right|}{\left|{{\mathbf{P}_{{{\mathbf{a}}_{k}}}}}\cup{{\mathbf{G}_{{{\mathbf{a}}_{k}}}}}\right|}-\frac{\left|\mathbf{C}\right|-\left|{{\mathbf{P}_{{{\mathbf{a}}_{k}}}}}\cup{{\mathbf{G}_{{{\mathbf{a}}_{k}}}}}\right|}{\left|\mathbf{C}\right|},\\ \mathrm{MGIoU}^{3D}=\frac{1}{3}\sum_{k=1}^{3}\mathrm{GIoU}_{k}^{1D}.\end{gathered}$ (5)
@@ 3.4 Total Loss and Training Strategy
Although SPAN promotes joint optimization of the 3D attributes and markedly improves both spatial and projection consistency, it can destabilize early training. Specifically, noisy initial predictions of the 3D box parameters can invalidate the assumptions underlying these high‑order constraints. To mitigate this, we adopt and adapt the Hierarchical Task Learning (HTL) strategy [22]. HTL applies a time‑varying weight ${\omega_{i}}\left(t\right)$ to each loss term, dynamically adjusting these weights at each epoch based on the learning progress of their prerequisite tasks.
@@
From a high-level perspective, as illustrated in Fig. 4, the overall learning procedure is organized into four sequential stages. Stage 1 focuses on 2D detection, covering object classification, 2D box localization, and projected center regression. Stage 2 addresses 3D dimension and rotation angle regression. Following MonoDGP [30], whose decoder is decoupled and initializes 3D object queries from 2D cues, we treat the tasks of Stage 1 as pre-tasks for Stage 2. Stage 3 addresses depth estimation. As this process relies on the geometric relationship between the 2D box attributes from Stage 1 and the 3D attributes from Stage 2, it is logically dependent on the outputs of both. Finally, Stage 4 introduces the Spatial-Projection Alignment, which depends on all preceding 3D attribute regression tasks. This staged design ensures that each task commences only after its prerequisite tasks have stabilized, thereby preserving training stability throughout the optimization process. The implementation of HTL is provided in the appendix Sec. C.
@@ A Detailed Discussion on Projection Constraint
In summary, the extrema of $f\left({{x_{n}},{z_{n}}}\right)$ over the point set $S$ are necessarily attained in at least one of the four line segments $(x_{\min},y_{n},z_{\min})$, $(x_{\min},y_{n},z_{\max})$, $(x_{\max},y_{n},z_{\min})$, and $(x_{\max},y_{n},z_{\max})$. Since the vertices of the 3D bounding box are precisely the endpoints of these segments, the maximum and minimum values of $u_{n}$ over $S$ are realized at these vertices, the identical argument holds for $v_{n}$.
@@
${u_{\min}}\leqslant{u_{n}}\leqslant{u_{\max}},{v_{\min}}\leqslant{v_{n}}\leqslant{v_{\max}}.$ (16)
@@ B Detailed Loss Function
$\begin{gathered}\mathcal{L}_{region}^{i}=1-\frac{2\cdot\sum_{j=1}^{N_{i}}p_{j}\cdot g_{j}}{\sum_{j=1}^{N_{i}}p_{j}+\sum_{j=1}^{N_{i}}g_{j}},\\ \mathcal{L}_{region}=\sum_{i=1}^{r}{\mathcal{L}_{region}^{i}},\end{gathered}$ (21)
@@ D Experiments on Waymo Open Dataset
E Compared with Prior Geometric Methods
While sharing geometric roots, SPAN innovates by formulating consistency as a differentiable, single-stage optimization objective, differing from prior works [25, 26]. The results of the comparative experiment are shown in Tab. 8.
Soft Constraints vs. Hard Solving.  Deep3DBox [25] uses rigid algebraic solving, where 2D noise propagates to catastrophic 3D failure (-0.81%). SPAN’s differentiable soft losses enable the network to learn resilience to 2D jitter, ensuring robustness.
Gradient Refinement vs. Active Regression.  Shift R-CNN [26] relies on heavy multi-stage Active Regression (+15ms). SPAN achieves refinement via end-to-end gradient optimization, yielding geometric gains in a single-stage pipeline at zero inference cost.
Geometric Strategy Mechanism Inference Cost (Batch Size of 1) $AP_{3D}$ (Mod.)
Baseline (MonoDGP [30]) Pure Regression 42ms (single RTX 3090) 22.34
+ Deep3DBox [25] Hard Algebraic Solving +5ms 21.53 (-0.81)
+ Shift R-CNN [26] Multi-stage Active Reg. +15ms 22.85 (+0.51)
+ SPAN (Ours) Soft Loss + HTL +0ms 23.26 (+0.92)
Table 8: Comparison of Geometric Strategies on KITTI val set.
F Why MGIoU?
To address questions on MGIoU [10] choice and whether gains are solely from HTL, we conducted a comprehensive ablation. As shown in Tab. 9. MGIoU provides superior optimization dynamics compared to baselines. (i) Unlike L1 constraint which treats corners independently, MGIoU enforces structural volumetric integrity. (ii) Unlike Exact IoU, MGIoU utilizes projection-based approximation to provide non-vanishing gradients for disjoint boxes, ensuring superior convergence. MGIoU outperforms L1 (+0.21%) and Exact IoU (+0.14%). By rigidly coupling depth and dimensions, it effectively mitigates depth bias by preventing the network from exploiting 2D projection ambiguities.
Method Config HTL $AP_{3D}$ (Mod.) $\Delta$
Baseline (MonoDGP) - 22.34 -
Baseline + HTL Only 22.56 +0.22
Baseline + HTL + $\mathcal{L}_{corner}$ (L1) 22.68 +0.34
Baseline + HTL + $\mathcal{L}_{corner}$ (Exact 3D IoU) 22.75 +0.41
Baseline + HTL + $\mathcal{L}_{corner}$ (MGIoU) 22.89 +0.55
Table 9: MGIoU performance analysis.
G Disentangling HTL & Geometric Alignment Gains
As shown in Tab. 10, we conducted an experiment replacing HTL with a simple Linear Weight Schedule to isolate gains. By using a naive Linear Weight Schedule to mitigate initial noise, SPAN achieves 22.95% AP, outperforming the HTL baseline. This confirms that gains stem primarily from the geometric losses. HTL further boosts performance (+0.31%) by adapting to training dynamics, acting as an amplifier for SPAN.
Method Config Schedule $AP_{3D}$ (Mod.) $\Delta$
Baseline None 22.34 -
Baseline + SPAN None 21.85 -0.49
Baseline HTL 22.56 +0.22
Baseline + SPAN Linear Weight Schedule 22.95 +0.61
Baseline + SPAN HTL 23.26 +0.92
Table 10: Decoupling Schedule vs. Geometric Gains.
H Interplay between Projection Alignment & Depth Bias
We analyze the interaction between Projection Alignment (PA) and depth bias.
PA mitigates Depth Bias. To address the inherent scale ambiguity ($z\propto f\cdot H_{3d}/h_{2d}$), PA utilizes reliable 2D boxes as geometric anchors, rectifying depth errors that pure regression permits. Tab. 11 validates that PA reduces long-range Depth MAE, directly contributing to $AP_{3D}$ gains.
Depth Loss with Geometry Prior regularizes PA. Conversely, depth bias loss by MonoDGP [30] prevents PA from converging to degenerate solutions. By penalizing projection violations, it ensures the geometric optimization remains physically consistent.
Method Depth MAE (m) $\downarrow$ by Distance $AP_{3D}$ (Mod.)
0-20m 20-40m 40m-$\infty$
MonoDGP 0.34 0.96 1.55 22.34
+ Proj. Align + HTL 0.36 (+0.02) 0.92 (-0.04) 1.50 (-0.05) 22.97
Table 11: Impact of Projection Alignment on Depth Bias.
I More Qualitative Results